# Beyond Slater’s Condition in Online CMDPs with Stochastic and Adversarial Constraints

### v1

@@ B.8 Positive Violation Bound
$=\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau,i}+\widehat{g}_{\tau,i})^{\top}q_{\tau}\right]^{+}$
$=\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau,i})^{\top}q_{\tau}+\widehat{g}_{\tau,i}^{\top}q_{\tau}\right]^{+}$
$=\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau,i})^{\top}q_{\tau}+\widehat{g}_{\tau,i}^{\top}q_{\tau}-\widehat{g}_{\tau,i}^{\top}\widehat{q}_{\tau}+\widehat{g}_{\tau,i}^{\top}\widehat{q}_{\tau}\right]^{+}$
$\leq\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau,i})^{\top}q_{\tau}+\widehat{g}_{\tau,i}^{\top}q_{\tau}-\widehat{g}_{\tau,i}^{\top}\widehat{q}_{\tau}+b_{\tau}^{\top}\widehat{q}_{\tau}\right]^{+}$ (34)
$\leq\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau,i})^{\top}q_{\tau}+b_{\tau}^{\top}\widehat{q}_{\tau}\right]^{+}+\|q_{\tau}-\widehat{q}_{\tau}\|_{1}$
$\leq\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau,i})^{\top}q_{\tau}+b_{\tau}^{\top}\widehat{q}_{\tau}\right]^{+}+2L|X|\sqrt{2t\ln\frac{2L}{\delta}}+3L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}}$ (35)
$\leq\sum_{\tau=1}^{t}\left[b_{\tau}^{\top}q_{\tau}+b_{\tau}^{\top}\widehat{q}_{\tau}\right]^{+}+5L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}},$ (36)
where Inequality (34) holds since $\widehat{q}_{\tau}\in\widehat{\Delta}_{t}(\mathcal{P}_{t})$, Inequality (35) follows from Lemma B.3 of [Rosenberg and Mansour, 2019b] with probability at least $1-2\delta$ and Inequality (36) holds by Lemma 4.1 with probability $1-11\delta$ jointly for each $i$ and $t$.
@@ Appendix C Technical Lemmas
$\left|\frac{1}{N_{t}(x,a)}\sum_{\tau\in\mathcal{T}_{t-1,x,a}}g_{\tau,i}(x,a)-\bar{g}_{i}(x,a)\right|\leq\sqrt{\frac{2\ln\frac{2m|X||A|T}{\delta}}{N_{t}(x,a)}}.$

### v2

@@ B.8 Positive Violation Bound
$=\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau-1,i}+\widehat{g}_{\tau-1,i})^{\top}q_{\tau}\right]^{+}$
$=\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau-1,i})^{\top}q_{\tau}+\widehat{g}_{\tau-1,i}^{\top}q_{\tau}\right]^{+}$
$=\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau-1,i})^{\top}q_{\tau}+\widehat{g}_{\tau-1,i}^{\top}q_{\tau}-\widehat{g}_{\tau-1,i}^{\top}\widehat{q}_{\tau}+\widehat{g}_{\tau-1,i}^{\top}\widehat{q}_{\tau}\right]^{+}$
$\leq\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau-1,i})^{\top}q_{\tau}+\widehat{g}_{\tau-1,i}^{\top}q_{\tau}-\widehat{g}_{\tau-1,i}^{\top}\widehat{q}_{\tau}+b_{\tau-1}^{\top}\widehat{q}_{\tau}\right]^{+}$ (31)
$\leq\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau-1,i})^{\top}q_{\tau}+b_{\tau-1}^{\top}\widehat{q}_{\tau}\right]^{+}+\|q_{\tau}-\widehat{q}_{\tau}\|_{1}$
$\leq\sum_{\tau=1}^{t}\left[(\bar{g}_{i}-\widehat{g}_{\tau-1,i})^{\top}q_{\tau}+b_{\tau-1}^{\top}\widehat{q}_{\tau}\right]^{+}+2L|X|\sqrt{2t\ln\frac{2L}{\delta}}+3L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}}$ (32)
$\leq\sum_{\tau=1}^{t}\left[b_{\tau-1}^{\top}q_{\tau}+b_{\tau-1}^{\top}\widehat{q}_{\tau}\right]^{+}+5L|X|\sqrt{2t|A|\ln\frac{2T|X||A|}{\delta}},$ (33)
where Inequality (31) holds since $\widehat{q}_{\tau+1}\in\widehat{\Delta}_{t}(\mathcal{P}_{t})$, Inequality (32) follows from Lemma B.3 of [Rosenberg and Mansour, 2019b] with probability at least $1-2\delta$ and Inequality (33) holds by Lemma 4.1 with probability $1-11\delta$ jointly for each $i$ and $t$.
@@ Appendix C Technical Lemmas
$\left|\frac{1}{N_{t}(x,a)}\sum_{\tau\in\mathcal{T}_{t,x,a}}g_{\tau,i}(x,a)-\bar{g}_{i}(x,a)\right|\leq\sqrt{\frac{2\ln\frac{2m|X||A|T}{\delta}}{N_{t}(x,a)}}.$